OpenStudy (anonymous):

What is 7 · 2 3/7? please help ILL FAN AND MEDAL

TheSmartOne (thesmartone):

First make 2 3/7 into a improper fraction

OpenStudy (anonymous):

how

TheSmartOne (thesmartone):

\(\sf\Large a \frac{b}{c} = \frac{ac+b}{c}\)

OpenStudy (anonymous):

ok so 7 x 3 + 2?

OpenStudy (anonymous):

23/7

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

is it 23/7? @TheSmartOne

TheSmartOne (thesmartone):

in my example, 2 would be 'a' 3 would be 'b' and 7 would be 'c' so try again :)

OpenStudy (anonymous):

7 x 2 + 3?

OpenStudy (anonymous):

17/ 7

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

correct now here's another hint: \(\sf\Large a \times \frac{b}{a} = b\) So what is \(\sf\Large 7 \times \frac{17}{7} = ?\) @midnighti

OpenStudy (anonymous):

is the fraction a division problem? like 17 divided by 7? and then I multiply by 7? @TheSmartOne

TheSmartOne (thesmartone):

look at my hint you won't need to divide you need to simplify also no need to tag me each time ._.

OpenStudy (anonymous):

I don't have much time for this I'm still only on question 1 in math and I just need the answer

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne correct now here's another hint: \(\sf\Large a \times \frac{b}{a} = b\) So what is \(\sf\Large 7 \times \frac{17}{7} = ?\) @midnighti \(\color{blue}{\text{End of Quote}}\)

TheSmartOne (thesmartone):

in your case, a = 7 and b = 17 do you understand? :P

OpenStudy (anonymous):

I never will understand fractions I'm stupid with them xD

TheSmartOne (thesmartone):

now you are making yourself believe that so you have an excuse to not learn